OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 90 yards and a height of 20 yards?

OpenStudy (callisto):

volume =1/3 x base area x height =1/3 x r^2 x h = 1/3 x 90^2 x 20 = 54000

OpenStudy (anonymous):

Why do you square the base height?

OpenStudy (callisto):

area of square = length x length =( length )^2

OpenStudy (anonymous):

Oh, okay! Thanks so much!

OpenStudy (callisto):

welcome :)